Determine the molarity of each of the following solutions:

a. 2.34 mol KCl in 2.300 L of solution

b. 0.56 g H2SO4 in 2 L of solution

c. 20.45 g Al(NO3)3 in 2000 mL of solution


1
Expert's answer
2022-02-23T22:06:02-0500

a) molarity "=\frac{2.34}{2.300}=1.017M"


b) molarity "=\frac{0.56}{98×2}=0.0029M"


c) molarity "=\frac{20.45}{223×2}=" 0.046M